# How to build the best medical image segmentation algorithm using foundation models: a comprehensive empirical study with Segment Anything Model

### v1

@@ Abstract
Automated segmentation is a fundamental medical image analysis task, which enjoys significant advances due to the advent of deep learning. While foundation models have been useful in natural language processing and some vision tasks for some time, the foundation model developed with image segmentation in mind - Segment Anything Model (SAM) - has been developed only recently and has shown similar promise. However, there are still no systematic analyses or “best-practice” guidelines for optimal fine-tuning of SAM for medical image segmentation. This work summarizes existing fine-tuning strategies with various backbone architectures, model components, and fine-tuning algorithms across 18 combinations, and evaluates them on 17 datasets covering all common radiology modalities. Our study reveals that (1) fine-tuning SAM leads to slightly better performance than previous segmentation methods, (2) fine-tuning strategies that use parameter-efficient learning in both the encoder and decoder are superior to other strategies, (3) network architecture has a small impact on final performance, (4) further training SAM with self-supervised learning can improve final model performance. We also demonstrate the ineffectiveness of some methods popular in the literature and further expand our experiments into few-shot and prompt-based settings. Lastly, we released our code and MRI-specific fine-tuned weights, which consistently obtained superior performance over the original SAM, at https://github.com/mazurowski-lab/finetune-SAM.
@@ 1 Introduction
Medical image segmentation refers to identifying the region of interest, typically anatomical structures or tumors, in an image. This process is essential in the field of medical image analysis, significantly contributing to the effectiveness of computer-aided diagnosis [1]. The introduction of deep learning techniques has led to a variety of automated segmentation methods [57, 85], further enhancing the precision and efficiency of analyzing medical images. Despite the progress made, it is still difficult to determine the most effective method for developing an optimal segmentation model, especially with the recent introduction of vision foundation models dedicated to segmentation. Therefore, in our study, we aim to explore and define the optimal strategies for developing automated segmentation algorithms specifically designed for medical images and with a focus on foundation models.
In the era of deep learning, a common practice for automated medical image segmentation is to directly train UNet [70] or its variants [75, 83] for a dataset and task. Recently, there has been a growing interest in the field of customizing vision foundation models towards specific segmentation tasks [96, 29]. Foundation models are (typically transformer-based [82]) neural networks pre-trained with extremely large datasets, such that they can generalize to a wide range of tasks and datasets. The seminal Segment Anything Model (SAM) [45] was proposed as the first large model designed with segmentation tasks in mind. SAM employs a prompt-based training method, aiming to generate target object’s segmentation masks in response to user-provided prompts, e.g. points or bounding boxes. It has demonstrated comparable or even superior zero-shot performance to fully supervised models in processing natural images. However, SAM’s performance is less impressive when directly applied to medical images [61, 21, 39], and the requirement of prompts makes the direct usage of it with medical imaging segmentation tasks challenging.
@@ 4.3.2 Results
Figure 8 presents the results of using SSLSAM as the weights before task-specific supervised training. We observe an improvement in the average performance of all MRIs across all architecture combinations besides En/Decoder + LoRA. Noticeably, when selecting En/Decoder + Adapter as the fine-tuning strategy, we observe consistent improvements over all MRI datasets, and the model achieves an average performance of , which surpasses the leading ViT-H based method by and the best UNet-based method (nnUNet) by . Contradictory to task-expansive prompt-based learning, the features learned during SSL are task agnostic, and since we only pre-train the encoder, there is no dependence on the task condition, which we have shown to hurt the performance. In the non-MRI datasets base, we observe a minor decrease in the performance. This can be due to the specificity of the MRI.
@@ 4.4.1 Implementation details
One of the popular scenarios for adopting fine-tuning strategies is when the target dataset has a limited number of annotated images, known as few-shot fine-tuning. Few-shot fine-tuning gets even more attention in the field of medical imaging segmentation, given the difficulty in securing a substantial quantity of expert annotations. We explore the effectiveness of different fine-tuning procedures under the following setting.
Our study primarily investigates 5-shot learning, selecting five images and their corresponding annotations randomly from the training set and validation set respectively. In the case of datasets derived from 3D volumes, these samples were chosen from distinct volumes to ensure diversity. In addition to the constraint of using only 5 training/validation images, we also modified the training hyper-parameters better suited for the few-shot setting. Specifically, we set a maximum of 1000 epochs for training, a batch size of 2, and did not include any early stopping criteria.
In this setting, we also investigated the previous three setups within a few-shot framework, including (1) task-specific fine-tuning exclusively, and (2) two pre-trained setups: task-expansive prompt-based learning and task-agnostic self-supervised learning, each followed by additional task-specific fine-tuning. To note, we cannot utilize the weights learned from task-expansive prompt-based learning since its training pipeline requires all labeled images. Thus, we consider the weights from MedSAM as an alternative. In this section, we investigate three key questions: (1) whether additional pre-training offers superior initial weights for the few-shot setting, (2) which fine-tuning strategy proves most effective in a few-shot context, (3) whether fine-tuning foundation models still work better than UNets in the few-shot setting?

### v2

@@ Abstract
Automated segmentation is a fundamental medical image analysis task, which enjoys significant advances due to the advent of deep learning. While foundation models have been useful in natural language processing and some vision tasks for some time, the foundation model developed with image segmentation in mind - Segment Anything Model (SAM) - has been developed only recently and has shown similar promise. However, there are still no systematic analyses or “best-practice” guidelines for optimal fine-tuning of SAM for medical image segmentation. This work summarizes existing fine-tuning strategies with various backbone architectures, model components, and fine-tuning algorithms across 18 combinations and evaluates them on 17 datasets covering all common radiology modalities. Our study reveals that (1) fine-tuning SAM leads to slightly better performance than previous segmentation methods, (2) fine-tuning strategies that use parameter-efficient learning in both the encoder and decoder are superior to other strategies, (3) network architecture has a small impact on final performance, and (4) further training SAM with self-supervised learning can improve final model performance. We also demonstrate the ineffectiveness of some methods popular in the literature and further expand our experiments into few-shot and prompt-based settings. Lastly, we released our code and MRI-specific fine-tuned weights, which consistently obtained superior performance over the original SAM, at https://github.com/mazurowski-lab/finetune-SAM.
@@ 1 Introduction
Medical image segmentation refers to identifying the region of interest, typically anatomical structures or tumors, in an image. This process is essential in the field of medical image analysis, contributing significantly to the effectiveness of computer-aided diagnosis [1]. The introduction of deep learning techniques has led to a variety of automated segmentation methods [57, 85], further enhancing the precision and efficiency of analyzing medical images. Despite the progress made, it is still difficult to determine the most effective method for developing an optimal segmentation model, especially with the recent introduction of vision foundation models dedicated to segmentation. Therefore, in our study, we aim to explore and define the optimal strategies for developing automated segmentation algorithms specifically designed for medical images and with a focus on foundation models.
In the era of deep learning, a common practice for automated medical image segmentation is to directly train UNet [70] or its variants [75, 83] for a dataset and task. Recently, there has been a growing interest in the field of customizing vision foundation models towards specific segmentation tasks [96, 29]. Foundation models are (typically transformer-based [82]) neural networks pre-trained with extremely large datasets, such that they can generalize to a wide range of tasks and datasets. The seminal Segment Anything Model (SAM) [45] was proposed as the first large model designed with segmentation tasks in mind. SAM employs a prompt-based training method, aiming to generate target object’s segmentation masks in response to user-provided prompts, e.g. points or bounding boxes. It has demonstrated comparable or even superior zero-shot performance to fully supervised models in processing natural images. However, SAM performance is less impressive when applied directly to medical images [61, 21, 39], and the requirement of prompts makes its direct usage with medical imaging segmentation tasks challenging.
@@ 4.3.2 Results
Figure 8 presents the results of using SSLSAM as the weights before task-specific supervised training. We observe an improvement in the average performance of all MRIs across all architecture combinations besides En/Decoder + LoRA. Noticeably, when selecting En/Decoder + Adapter as the fine-tuning strategy, we observe consistent improvements over all MRI datasets, and the model achieves an average performance of , which surpasses the leading ViT-H based method by and the best UNet-based method (nnUNet) by . Contradictory to task-expansive prompt-based learning, the features learned during SSL are task agnostic, and since we only pre-train the encoder, there is no dependence on the task condition, which we have shown to hurt the performance. In the non-MRI data sets base, we observe a minor decrease in performance. This may be due to the specificity of the MRI.
@@ 4.4.1 Implementation details
One of the popular scenarios for adopting fine-tuning strategies is when the target dataset has a limited number of annotated images, known as few-shot fine-tuning. Few-shot fine-tuning gets even more attention in the field of medical imaging segmentation, given the difficulty in securing a substantial quantity of expert annotations. We explore the effectiveness of different fine-tuning procedures in the following setting.
Our study primarily investigates 5-shot learning, selecting five images and their corresponding annotations randomly from the training set and validation set, respectively. In the case of datasets derived from 3D volumes, these samples were chosen from distinct volumes to ensure diversity. In addition to the constraint of using only 5 training/validation images, we also modified the training hyper-parameters better suited for the few-shot setting. Specifically, we set a maximum of 1000 epochs for training, a batch size of 2, and did not include any early stopping criteria.
In this setting, we also investigated the previous three setups within a few-shot framework, including (1) task-specific fine-tuning exclusively, and (2) two pre-trained setups: task-expansive prompt-based learning and task-agnostic self-supervised learning, each followed by additional task-specific fine-tuning. Notably, we cannot use the weights learned from task-expansive prompt-based learning as its training pipeline requires all labeled images. Thus, we consider the MedSAM weights as an alternative. In this section, we investigate three key questions: (1) whether additional pre-training offers superior initial weights for the few-shot setting, (2) which fine-tuning strategy proves most effective in a few-shot context, and (3) whether fine-tuning foundation models still work better than UNets in the few-shot setting?